OpenStudy (anonymous):

4(a+3)-5=3a+8 solve for a please help.

OpenStudy (anonymous):

For the first part \[4(a+3)=?\]

OpenStudy (anonymous):

\[ 4(a+3)=4a+4*3 = 4a + 12\] So now \[4a+12-5=3a+8\] If you get the a's all on the left, and all the numbers on the right, then problem solved! ^_^

OpenStudy (anonymous):

can you get it from there?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

What'd ya get?